OpenStudy (anonymous):

please help? medals awarded! Multiply and simplify: x-6/x^2+4x-32*x-4/x+2

OpenStudy (anonymous):

can you please type it in equation mode

OpenStudy (anonymous):

it's hard to understand

OpenStudy (campbell_st):

factorise the quadratic in the 1st fraction \[(x -6)/(x + 8)(x - 4)\] the problem is now \[((x - 6) \times ( x - 4))/((x + 8)\times(x - 4)\times(x + 2))\] there is a common factor of (x -4) that can be cancelled \[(x-6)/((x + 8)(x + 2))\] expand the denominator which gives the final answer \[(x - 6)/(x^2 + 10x + 16)\]

OpenStudy (anonymous):

a. x-6/x+2 b.x-6/(x+8)(x+2) c.x-3/x8 which one of these?

OpenStudy (campbell_st):

the line before my final answer contains answer b

OpenStudy (anonymous):

ok thankyou!